OpenStudy (anonymous):

An object is falling from a height of 7.5 meters. At what height will its velocity be 7 meters/second? A. 49 meters B. 7.5 meters C. 7.0 meters D. 5.0 meters E. 2.5 meters

OpenStudy (anonymous):

according the formula \[v ^{2}=v _{0}^{2}+2as\] \[v=7 m/s\] \[v _{0}=0 m/s\] \[a=9.8 m/s^{2}\] the value of \[s=2.5m\] then \[7.5m-2.5m=5m\] so the answer is D

OpenStudy (anonymous):

Apply Equation of Motion. \[V^{2} = U^{2} + 2as\] (7)^2 = 2 * g * s1 g = gravitational constant s1 = 49/2g Distance from the ground = 7.5- 49/2g =